# Supervised heterogeneous network estimation via survival-based Bayesian graphical models

### v2

@@ 2.2 Prior specification
$\omega_{k,jl}\mid\gamma_{k,jl}\sim\gamma_{k,jl}\operatorname{LP}\left(\omega_{k,jl};v_{1}\right)+\left(1-\gamma_{k,jl}\right)\operatorname{LP}\left(\omega_{k,jl};v_{0}\right),\gamma_{k,jl}\sim\operatorname{Bern}\left(p_{1}\right),\text{ for }j<l,$
$\beta_{k,j}\sim\operatorname{LP}\left(\beta_{k,j};1/\lambda_{1}\right),\mu_{k,j}\sim\operatorname{LP}\left(\mu_{k,j};1/\lambda_{2}\right),g_{i}\mid\bm{\pi_{g}}\sim\operatorname{Multi}\left(\pi_{1},\pi_{2},\ldots,\pi_{K}\right).$
@@
Here, $\operatorname{Exp}(\omega_{k,jj};\tau_{0})$ is the Exponential distribution with parameter $\tau_{0}$, $\gamma_{k,jl}$ is a binary indicator corresponding to $\omega_{k,jl}$, $\operatorname{LP}(y;v)$ is the Laplace distribution with a scale parameter $v$, $p_{1}\in(0,1)$ is the prior probability for the Bernoulli distribution, and $\bm{\pi}_{g}=\left(\pi_{1},\ldots,\pi_{K}\right)^{\top}$ is the vector composed of the Multinomial distribution’s parameters, which specifies the prior probabilities that each sample is drawn from different subgroups. $\tau_{0}$, $\lambda_{1}$, $\lambda_{2}$, and $v_{1}>v_{0}$ are the positive tuning parameters.
@@
In the proposed Bayesian framework, we first introduce a weakly informative Exponential prior to the diagonal elements $\omega_{k,jj}$’s of the precision matrices, which are not subject to shrinkage, following [10]. We introduce a spike-and-slab Laplace prior [23, 28] for $\omega_{k,jl}$ based on the indicator $\gamma_{k,jl}$, for identifying individual sparse patterns of the $K$ networks. With this prior, when $\gamma_{k,jl}=1$, the corresponding entry $\omega_{k,jl}$ comes from the slab component with a relative large variance $2v_{1}^{2}$ and is likely to take a value away from zero; when $\gamma_{k,jl}=0$, $\omega_{k,jl}$ comes from the spike component with a small variance $2v_{0}^{2}$ that can induce values towards zero. Compared to the Laplace prior, the spike-and-slab Laplace prior involves two different scale parameters $v_{1}$ and $v_{0}$ to distinguish active from ignorable coefficients and can nonlinearly adapt to coefficients with $\gamma_{k,jl}$, achieving “selective shrinkage” with a smaller bias. This Laplace distribution-based prior can automatically achieve sparsity via the $L_{1}$ norm on $\omega_{k,jl}$. This advances from the commonly adopted spike-and-slab Gaussian prior which yields nonsparse posterior modes for $\omega_{k,jl}$’s and relies on a post-data thresholding-based selection strategy utilizing $\gamma_{k,jl}$. The strategy that is based on the estimation and sparsity of the continuous $\omega_{k,jl}$ can lead to more effective identification of sparse posterior patterns [23] and can also facilitate a simpler theoretical study and a more efficient computation.
Additionally, to accommodate the sparse structure of the truly associated predictors and predictor means, sparse priors are assumed for ${\beta}_{k,j}$’s and $\mu_{k,j}$’s for regularized estimation. As the estimation of ${\beta}_{k,j}$’s and $\mu_{k,j}$’s is less challenging than $\omega_{k,jl}$, we choose the relatively simple Laplace prior for both the regression coefficients and predictor means, which involves fewer parameters and less computational complexity.
@@ 2.3 Additional priors for accommodating common structures
To borrow information from cross-subgroup similarity, we further introduce a latent vector $\bm{\theta}_{jl}$ for $\bm{\omega}_{jl}$ and assume a Gaussian distribution with the precision matrix based on a similarity-based matrix $\bm{L}^{(jl)}$,
$\bm{\theta}_{jl}\sim\mathcal{N}\left(\bm{0},\left(u\bm{L}^{(jl)}\right)^{-1}\right),\bm{\theta}_{jl}\mid\bm{\omega}_{jl}\sim{I}\left\{\bm{\theta}_{jl}=\bm{\omega}_{jl}\right\},$
with $u$ being the positive tuning parameter. Here, $\bm{\theta}_{jl}\sim\mathcal{N}\left(\bm{0},\left(u\bm{L}^{(jl)}\right)^{-1}\right)$ can be interpreted as a generative model in which a zero vector is observed from a Gaussian distribution parameterized by $\bm{\theta}_{jl}$, i.e., $f\left(\bm{0}\mid\bm{\theta}_{jl}\right)=\sqrt{u\left|\bm{L}^{(jl)}\right|/(2\pi)^{K}}\exp\left\{-u\left(\bm{0}-\bm{\theta}_{jl}\right)^{\top}\bm{L}^{(jl)}\left(\bm{0}-\bm{\theta}_{jl}\right)/2\right\}$. This interpretation follows the hierarchical framework of [30], where priors are constructed through pseudo-observations. We define $L^{(jl)}_{k,k}=(K-1)/\left(\theta_{k,jl}^{2}+\epsilon^{2}\right)$ and $L^{(jl)}_{k,k^{\prime}}=-1/\sqrt{\left(\theta_{k,jl}^{2}+\epsilon^{2}\right)\left(\theta_{k^{\prime},jl}^{2}+\epsilon^{2}\right)}$ for $1\leq k\neq k^{\prime}\leq K$ and $1\leq j<l\leq p$, with $\epsilon$ being a small constant. Thus, for each $\bm{\theta}_{jl}=(\theta_{1,jl},\ldots,\theta_{K,jl})^{\top}$, the prior involves $\bm{\theta}_{jl}^{\top}\bm{L}^{(jl)}\bm{\theta}_{jl}=0.5\sum_{k\neq k^{\prime}}\left(\theta_{k,jl}/\sqrt{\theta_{k,jl}^{2}+\epsilon^{2}}-\theta_{k^{\prime},jl}/\sqrt{\theta_{k^{\prime},jl}^{2}+\epsilon^{2}}\right)^{2}$, which is an approximation to the Bayesian counterpart of the sign-based constraint penalty, that is $0.5\sum_{k\neq k^{\prime}}\left[\operatorname{sgn}(\theta_{k,jl})-\operatorname{sgn}(\theta_{k^{\prime},jl})\right]^{2}$. Combined with the indicator function ${I}\left\{\bm{\theta}_{jl}=\bm{\omega}_{jl}\right\}$, the proposed approach encourages the pairwise relationships of predictors across subgroups to have similar signs, leading to networks with a shared pattern of sparsity as well as edges with shared positive or negative relationships. This significantly advances from the previous research based on the group penalty [7, 12, 14, 6], which utilizes group selection without appropriately accounting for the similarity of the estimated values and fails to sufficiently address the relationships between different networks [15]. This is also more flexible than the strategy based on fused lasso [7, 11, 29], which promotes both shared directions and magnitudes of edges. Similar prior assumptions can also be applied to $\bm{\beta}$ and $\bm{\mu}$ when considering the common structures within regression coefficients and predictors’ means.
@@ 2.4 Posterior Inference
$f\left(\bm{\beta},\bm{\mu},\bm{\Omega},\bm{\theta},\bm{\gamma},\bm{g}\mid\bm{t},\bm{\delta},\textbf{X};\bm{\tau},\bm{\pi_{g}}\right)\propto$ $f\left(\bm{t},\bm{\delta},\textbf{X}\mid\bm{\beta},\bm{\mu},\bm{\Omega},\bm{\tau},\bm{g}\right)f\left(\bm{\beta}\right)f\left(\bm{\mu}\right)f\left(\bm{\Omega}\mid\bm{\gamma}\right)$ (2)
×$\times f\left(\bm{\gamma}\right)f\left(\bm{\theta}\mid\bm{\Omega}\right)f\left(\bm{0}\mid\bm{\theta}\right)f\left(\bm{g}\mid\bm{\pi_{g}}\right),$
with $f\left(\bm{0}\mid\bm{\theta}\right)=\prod_{j=1}^{p}\prod_{l>j}f\left(\bm{0}\mid\bm{\theta}_{jl}\right)$. We note that since the distribution $f\left(\bm{\theta}\mid\bm{\Omega}\right)$ is the indicator function, there is no need to estimate $\bm{\theta}$ separately. Figure S1 in the supplementary materials gives a graphical representation of the Bayesian model.
Based on (2), we develop an Expectation Maximization (EM) algorithm for posterior inference of the parameter set $\bm{\Psi}=\operatorname{vec}\left(\bm{\beta},\bm{\tau},\bm{\mu},\bm{\Omega}\right)$. Sensitivity analyses on hyperparameters confirm that our approach is robust across a range of plausible settings. The Bayesian Information Criterion (BIC) is adopted to choose optimal values of $v_{0}$ and $u$. We refer to Sections S2 and S3 of the supplementary materials for details.
2.5 Theoretical properties
Denote $\bm{\Psi}^{*}=\operatorname{vec}\left(\bm{\beta}^{*},\bm{\tau}^{*},\bm{\mu}^{*},\bm{\Omega}^{*}\right)$ as the vector composed of the unknown true parameters. For the $k$th subgroup, denote $\mathcal{U}_{k}^{*}=\left\{j:\mu_{k,j}^{\ast}\neq 0\right\}$ and $\mathcal{V}_{k}^{*}=\left\{j:\beta_{k,j}^{\ast}\neq 0\right\}$ as the sets of the nonzero elements of the true predictor means and regression parameters, and $\mathcal{W}_{k}^{*}=\left\{(j,l):j\neq l,\omega_{k,jl}^{\ast}\neq 0\right\}$ as the set of the nonzero off-diagonal elements of the true precision matrix. For a set $\mathcal{S}$, denote $\left|\mathcal{S}\right|$ as its cardinality. Define the sparsity parameters as $d_{\bm{\mu}}=\left|\cup_{k=1}^{K}\mathcal{U}_{k}^{*}\right|,d_{\bm{\beta}}=\left|\cup_{k=1}^{K}\mathcal{V}_{k}^{*}\right|$, and $s=\left|\cup_{k=1}^{K}\mathcal{W}_{k}^{*}\right|$. Define $\Xi$ as some non-empty convex set of parameters $\bm{\Psi}$ in $\mathbb{R}^{Kp^{2}+2Kp+2K}$. Denote ${\rho}_{k,i}^{\bm{\Psi}}$ as the probability that the $i$th subject belongs to the $k$th subgroup given $\bm{{\Psi}}$.
Furthermore, for a $q$-dimensional vector $\bm{\nu}$, denote $\left\|\bm{\nu}\right\|_{\max}=\max_{1\leq j\leq q}\left|\nu_{j}\right|$. For a matrix $\bm{M}\in\mathbb{R}^{q_{1}\times q_{2}}$, define its minimum, maximum, and max-induced norms as $\|\bm{M}\|_{\min}=\min_{i,j}\left|M_{ij}\right|$, $\|\bm{M}\|_{\max}=\max_{i,j}\left|M_{ij}\right|$, and $\|\bm{M}\|_{\infty}=\max_{i=1,\ldots,q_{1}}\sum_{j=1}^{q_{2}}\left|M_{ij}\right|$, respectively. For a square matrix $\bm{A}\in\mathbb{R}^{p\times p}$, let $\lambda_{\min}(\bm{A})$ and $\lambda_{\max}(\bm{A})$ denote its smallest and largest eigenvalues. The following conditions are assumed.
Assumption 1
Suppose $\underline{C}_{1}\leq\min_{k}\lambda_{\min}\left(\bm{\Omega}_{k}^{*}\right)\leq\max_{k}\lambda_{\max}\left(\bm{\Omega}_{k}^{*}\right)\leq\bar{C}_{1}$, where $\underline{C}_{1}$ and $\bar{C}_{1}$ are positive constants. Moreover, we denote ¨$\ddot{\ell}_{1}\left(\bm{\Psi},\bar{\bm{\Psi}}\right)$ and ¨$\ddot{\ell}_{2}\left(\bm{\Psi},\bar{\bm{\Psi}}\right)$ as the second derivatives of $\sum_{i}\sum_{k}{\rho}_{k,i}^{\bm{\Psi}}\delta_{i}\log f_{Z}\left(t_{i}\mid\bm{x}_{i},\bar{\bm{\beta}}_{k};\bar{\beta}_{0k},\bar{\tau}_{k}\right)$ and $\sum_{i}\sum_{k}{\rho}_{k,i}^{\bm{\Psi}}(1-\delta_{i})\log S_{Z}\left(t_{i}\mid\bm{x}_{i},\bar{\bm{\beta}}_{k};\bar{\beta}_{0k},\bar{\tau}_{k}\right)$ with respect to $\operatorname{vec}\left({\bm{\beta}},\bm{\tau}\right)$, respectively. Then for any $\underline{\bm{\Psi}}\in\Xi$, $\bm{\Psi}\in\mathcal{B}_{\alpha}\left(\bm{\Psi}^{*}\right)=\left\{\bm{\Psi}\in\Xi:\left\|\bm{\Psi}-\bm{\Psi}^{*}\right\|_{2}\right.$ $\left.\leq\alpha\right\}$ and $\bar{\bm{\Psi}}\in\left\{\bar{\bm{\Psi}}\mid\left\|\bar{\bm{\Psi}}-\bm{\Psi}^{*}\right\|_{2}\leq 2\alpha\right\}$, there exists $\widetilde{\eta}>0$ satisfying $\lambda_{\max}\left(\ddot{\ell}_{1}\left(\bm{\Psi},\bar{\bm{\Psi}}\right)\right)\leq-\widetilde{\eta}$ and $\lambda_{\max}\left(\ddot{\ell}_{2}\left(\bm{\Psi},\underline{\bm{\Psi}}\right)\right)\leq 0.$
Assumption 2
@@ 2.5 Theoretical properties
Assumption 3
$\left\|\bm{\beta}^{*}\right\|_{\max}$, $\left\|\bm{\tau}^{\ast}\right\|_{\max}$,
$\left\|\bm{\mu}^{*}\right\|_{\max}$, and $\max_{k}\left\|\bm{\Omega}_{k}^{*}\right\|_{\infty}$ are bounded from above, and $\left\|\bm{\tau}^{\ast}\right\|_{\min}$ is bounded from below. Furthermore, $\max\limits_{i=1,\cdots,n}{c}_{i}\leq\bar{C}_{2}$ and $\max\limits_{i=1,\cdots,n}\left\|\bm{x}_{i}\right\|_{\max}\leq\bar{C}_{3}$ for some constants $\bar{C}_{2}$ and $\bar{C}_{3}$.
Assumption 4
The number of subgroups $K$ satisfies $K\sqrt{\log n/p}=o\left(1\right)$.
Assumption 5
Assume the tuning parameters satisfy $u=O\left(\sqrt{{n(Ks+p)\log p}/{(Kp^{2}s^{2})}}\right)$, $\tau_{0}=O\left(\sqrt{n{(Ks+p)K\log p}/p}\right)$, $1/v_{0}=O\left(\sqrt{nK(Ks+p)\log p/s}\right)$, $1/v_{1}=O\left(\sqrt{nK(Ks+p)\log p/s}\right)$, and $v_{0}<v_{1}$.
Assumption 1 is a restricted strong concave condition for the log-likelihood function [13, 2]. Assumption 2 is the sufficient separability condition, which requires that the probability of each subject belonging to a subgroup is either close to 1 or 0. Such an assumption is often assumed in the literature [12, 22]. Assumption 3 puts constraints on the bounds of the true parameters and the predictors, which are very common in the relevant literature. Assumption 4 allows $K$ to grow at a rate slower than $\sqrt{p/\log n}$, which is also considered in the related literature [17, 8]. Assumption 5 restricts the rate of the tuning parameters, which are commonly assumed in high-dimensional studies [28, 31].

### v4

@@ 2.2 Prior specification
$\omega_{k,jl}\mid\gamma_{k,jl}\sim\gamma_{k,jl}\operatorname{LP}\left(\omega_{k,jl};v_{1}\right)+\left(1-\gamma_{k,jl}\right)\operatorname{LP}\left(\omega_{k,jl};v_{0}\right),\text{ for }j<l,$
$\gamma_{k,jl}\sim\operatorname{Bern}\left(p_{1}\right),\text{ for }j<l,$
$\beta_{k,j}\sim\operatorname{LP}\left(\beta_{k,j};1/\lambda_{1}\right),$
$\mu_{k,j}\sim\operatorname{LP}\left(\mu_{k,j};1/\lambda_{2}\right),$
$g_{i}\mid\boldsymbol{\pi_{g}}\sim\operatorname{Multi}\left(\pi_{1},\pi_{2},\ldots,\pi_{K}\right).$
@@
Here, $\operatorname{Exp}(\omega_{k,jj};\tau_{0})$ is the Exponential distribution with parameter $\tau_{0}$, $\gamma_{k,jl}$ is a binary indicator corresponding to $\omega_{k,jl}$, $\operatorname{LP}(y;v)$ is the Laplace distribution with a scale parameter $v$, $p_{1}\in(0,1)$ is the prior probability for the Bernoulli distribution, and $\boldsymbol{\pi}_{g}=\left(\pi_{1},\ldots,\pi_{K}\right)^{\top}$ is the vector composed of the Multinomial distribution’s parameters, which specifies the prior probabilities that each sample is drawn from different subgroups. $\tau_{0}$, $\lambda_{1}$, $\lambda_{2}$, and $v_{1}>v_{0}$ are the positive tuning parameters.
@@
In the proposed Bayesian framework, we first introduce a weakly informative Exponential prior to the diagonal elements $\omega_{k,jj}$’s of the precision matrices, which are not subject to shrinkage, following Gan et al. (2019). We introduce a spike-and-slab Laplace prior (Ročková, 2018; Yang et al., 2021) for $\omega_{k,jl}$ based on the indicator $\gamma_{k,jl}$, for identifying individual sparse patterns of the $K$ networks. With this prior, when $\gamma_{k,jl}=1$, the corresponding entry $\omega_{k,jl}$ comes from the slab component with a relative large variance $2v_{1}^{2}$ and is likely to take a value away from zero; when $\gamma_{k,jl}=0$, $\omega_{k,jl}$ comes from the spike component with a small variance $2v_{0}^{2}$ that can induce values towards zero. Compared to the Laplace prior, the spike-and-slab Laplace prior involves two different scale parameters $v_{1}$ and $v_{0}$ to distinguish active from ignorable coefficients and can nonlinearly adapt to coefficients with $\gamma_{k,jl}$, achieving “selective shrinkage” with a smaller bias. The spike-and-slab Laplace prior utilizes latent indicators $\gamma_{k,jl}$ to facilitate a rigorous assessment of Bayesian uncertainty. Rather than relying on the point estimates of $\omega_{k,jl}$, this approach performs network construction via the posterior inclusion probabilities derived from $\gamma_{k,jl}$. This probabilistic framework effectively quantifies model-selection uncertainty while ensuring theoretical tractability and computational efficiency.
Additionally, to accommodate the sparse structure of truly associated predictors and predictor means, sparse priors are assumed for ${\beta}_{k,j}$ and $\mu_{k,j}$ for regularized estimation. Since only a subset of molecular means contributes to subgroup differentiation, imposing sparsity on $\mu_{k,j}$ is essential in high-dimensional clustering, and estimators without such regularization can be sub-optimal (Hao et al., 2018). Given that estimating ${\beta}_{k,j}$ and $\mu_{k,j}$ is generally less demanding than estimating the precision matrix entries $\omega_{k,jl}$, we employ a Laplace prior for both sets of parameters. This choice provides a suitable trade-off between effective regularization and computational efficiency.
@@ 2.3 Additional priors for accommodating common structures
To borrow information from cross-subgroup similarity, we further introduce a latent vector $\boldsymbol{\theta}_{jl}=\left(\theta_{1,jl},\ldots,\theta_{K,jl}\right)^{\top}$ for $\boldsymbol{\omega}_{jl}=\left(\omega_{1,jl},\ldots,\omega_{K,jl}\right)^{\top}$ and assume a Gaussian distribution with the precision matrix based on a similarity-based matrix $\boldsymbol{L}^{(jl)}$,
$\boldsymbol{\theta}_{jl}\sim\mathcal{N}\left(\boldsymbol{0},\left(u\boldsymbol{L}^{(jl)}\right)^{-1}\right),\quad\boldsymbol{\theta}_{jl}\mid\boldsymbol{\omega}_{jl}\sim{I}\left\{\boldsymbol{\theta}_{jl}=\boldsymbol{\omega}_{jl}\right\},$
with $u$ being the positive tuning parameter. Here, $\boldsymbol{\theta}_{jl}\sim\mathcal{N}\left(\boldsymbol{0},\left(u\boldsymbol{L}^{(jl)}\right)^{-1}\right)$ can be interpreted as a generative model in which a zero vector is observed from a Gaussian distribution parameterized by $\boldsymbol{\theta}_{jl}$, i.e., $f\left(\boldsymbol{0}\mid\boldsymbol{\theta}_{jl}\right)=\sqrt{u\left|\boldsymbol{L}^{(jl)}\right|/(2\pi)^{K}}\exp\left\{-u\left(\boldsymbol{0}-\boldsymbol{\theta}_{jl}\right)^{\top}\boldsymbol{L}^{(jl)}\left(\boldsymbol{0}-\boldsymbol{\theta}_{jl}\right)/2\right\}$. This interpretation follows the hierarchical framework of Zhe et al. (2013), where priors are constructed through pseudo-observations. We define $L^{(jl)}_{k,k}=(K-1)/\left(\theta_{k,jl}^{2}+\epsilon^{2}\right)$ and $L^{(jl)}_{k,k^{\prime}}=-1/\sqrt{\left(\theta_{k,jl}^{2}+\epsilon^{2}\right)\left(\theta_{k^{\prime},jl}^{2}+\epsilon^{2}\right)}$ for $1\leq k\neq k^{\prime}\leq K$ and $1\leq j<l\leq p$, with $\epsilon$ being a small constant. Thus, for each $\boldsymbol{\theta}_{jl}$, the prior involves $\boldsymbol{\theta}_{jl}^{\top}\boldsymbol{L}^{(jl)}\boldsymbol{\theta}_{jl}=0.5\sum_{k\neq k^{\prime}}\left(\theta_{k,jl}/\sqrt{\theta_{k,jl}^{2}+\epsilon^{2}}-\theta_{k^{\prime},jl}/\sqrt{\theta_{k^{\prime},jl}^{2}+\epsilon^{2}}\right)^{2}$, which is an approximation to the Bayesian counterpart of the sign-based constraint penalty $0.5\sum_{k\neq k^{\prime}}\left[\operatorname{sgn}(\theta_{k,jl})-\operatorname{sgn}(\theta_{k^{\prime},jl})\right]^{2}$. Combined with the indicator function ${I}\left\{\boldsymbol{\theta}_{jl}=\boldsymbol{\omega}_{jl}\right\}$, the proposed approach encourages the pairwise relationships of predictors across subgroups to have similar signs, leading to networks with a shared pattern of sparsity as well as edges with shared positive or negative relationships. This significantly advances from the previous research based on the group penalty (Danaher et al., 2014; Hao et al., 2018; Li et al., 2019; Dai and Jin, 2022), which utilizes group selection without appropriately accounting for the similarity of the estimated values and fails to sufficiently address the relationships between different networks. This is also more flexible than the strategy based on fused lasso (Danaher et al., 2014; Gao et al., 2016; Zhang et al., 2021), which promotes both shared directions and magnitudes of edges.
@@ 2.4 Posterior inference
$f\left(\boldsymbol{\beta},\boldsymbol{\mu},\boldsymbol{\Omega},\boldsymbol{\theta},\boldsymbol{\gamma},\boldsymbol{g}\mid\boldsymbol{t},\boldsymbol{\delta},\textbf{X};\boldsymbol{\tau},\boldsymbol{\pi_{g}}\right)\propto$ $f\left(\boldsymbol{t},\boldsymbol{\delta},\textbf{X}\mid\boldsymbol{\beta},\boldsymbol{\mu},\boldsymbol{\Omega},\boldsymbol{\tau},\boldsymbol{g}\right)f\left(\boldsymbol{\beta}\right)f\left(\boldsymbol{\mu}\right)f\left(\boldsymbol{\Omega}\mid\boldsymbol{\gamma}\right)f\left(\boldsymbol{\gamma}\right)$ (2)
$\times f\left(\boldsymbol{\theta}\mid\boldsymbol{\Omega}\right)f\left(\boldsymbol{0}\mid\boldsymbol{\theta}\right)f\left(\boldsymbol{g}\mid\boldsymbol{\pi_{g}}\right){\prod_{k=1}^{K}I\left\{\boldsymbol{\Omega}_{k}\succ 0\right\}},$
where $f\left(\boldsymbol{0}\mid\boldsymbol{\theta}\right)=\prod_{j=1}^{p}\prod_{l>j}f\left(\boldsymbol{0}\mid\boldsymbol{\theta}_{jl}\right)$ and $I\left\{\boldsymbol{\Omega}_{k}\succ 0\right\}$ ensures that the precision matrix $\boldsymbol{\Omega}_{k}$ is positive definite. We note that since the distribution $f\left(\boldsymbol{\theta}\mid\boldsymbol{\Omega}\right)$ is the indicator function, there is no need to estimate $\boldsymbol{\theta}$ separately. Figure S1 in the supplementary materials gives a graphical representation of the Bayesian model.
@@
In the second step, the uncertainty of the edge inclusion is quantified using the marginal MAP plug-in posterior inclusion probability (PIP): $\tilde{\gamma}_{k,jl}=\mathbb{P}(\gamma_{k,jl}=1\mid\tilde{\omega}_{k,jl})$. Finally, in the third step, the network sparsity pattern is determined by thresholding $\tilde{\gamma}_{k,jl}$, with entries exceeding a prespecified level such as 0.5 identified as edges. Further details are provided in Section S2 and Algorithm 1 of the supplementary materials.
Sensitivity analyses on hyperparameters confirm that our approach is robust across a range of plausible settings. The Bayesian Information Criterion (BIC) is adopted to choose optimal values of $v_{0}$ and $u$, which is defined as $2\log f\left(\boldsymbol{t},\boldsymbol{\delta},\mathbf{X}\mid\tilde{\boldsymbol{\Psi}},\hat{\boldsymbol{\pi}}_{g}\right)+\log n\hat{S}$ where $\tilde{\boldsymbol{\Psi}}$ and $\hat{\boldsymbol{\pi}}_{g}$ are the final estimates obtained from the proposed algorithm and $\hat{S}$ denotes the number of nonzero elements in the sparse parameter estimates (Hao et al., 2018; Ren et al., 2022). We refer to Sections S2 and S3 of the supplementary materials for details.
3 Theoretical properties
Denote the vector of unknown true parameters by $\boldsymbol{\Psi}^{\star}:=\left(\boldsymbol{\Theta}^{\star},\boldsymbol{\tau}^{\star}\right)=\operatorname{vec}\left(\boldsymbol{\beta}^{\star},\boldsymbol{\mu}^{\star},\boldsymbol{\Omega}^{\star},\boldsymbol{\tau}^{\star}\right)$. For the $k$th subgroup, let $\mathcal{U}_{k}^{\star}=\left\{j:\mu_{k,j}^{\ast}\neq 0\right\}$ and $\mathcal{V}_{k}^{\star}=\left\{j:\beta_{k,j}^{\ast}\neq 0\right\}$ denote the index sets of nonzero true predictor means and regression coefficients, respectively, and let $\mathcal{W}_{k}^{\star}=\left\{(j,l):j\neq l,\omega_{k,jl}^{\ast}\neq 0\right\}$ denote the set of nonzero off-diagonal elements of the true precision matrix. For a set $\mathcal{S}$, let $\left|\mathcal{S}\right|$ denote its cardinality. Define $d_{\boldsymbol{\mu}}=\left|\cup_{k=1}^{K}\mathcal{U}_{k}^{\star}\right|,d_{\boldsymbol{\beta}}=\left|\cup_{k=1}^{K}\mathcal{V}_{k}^{\star}\right|$, and $s=\left|\cup_{k=1}^{K}\mathcal{W}_{k}^{\star}\right|$. For a matrix $\boldsymbol{M}\in\mathbb{R}^{q_{1}\times q_{2}}$, define $\|\boldsymbol{M}\|_{\infty}=\max_{i=1,\ldots,q_{1}}\sum_{j=1}^{q_{2}}\left|M_{ij}\right|$.
In Section S4 of the supplementary materials, we specify the assumed conditions, including log-likelihood regularity, subgroup separability, parameter boundedness, and the relative growth rates of tuning parameters and $K$ alongside the minimal signal strength assumption. Similar conditions have been adopted in prior studies (Hao et al., 2021, 2018; Ren et al., 2022; Zhong et al., 2023).